OpenStudy (anonymous):

Simplify to lowest terms, if possible: (14 t^3)/(49 t^13)

OpenStudy (anonymous):

\[\huge \frac{(14 \times t^3)}{(49 \times t^{13})}\] \[\huge= \frac{(2 \times 7)}{(7 \times 7 \times t^{13-3})}\] \[\huge= \frac{(2 )}{(7 \times t^{10})} = \frac{2}{7t^{10}}\]

OpenStudy (anonymous):

your the best